OpenStudy (anonymous):

Judy and Kim have rectangular-shaped vegetable gardens. The width and area of Kim’s garden is double the width and area of Judy’s garden. Which statement is true? Kim’s garden has the same perimeter as Judy’s garden. The length of Kim’s garden is double the length of Judy’s garden. Kim’s garden has the same length as Judy’s garden. The length of Kim’s garden is four times the length of Judy’s garden.

OpenStudy (anonymous):

I believe its "Kim's garden has the same length as Judy's garden."

OpenStudy (anonymous):

Kims garden is the same length as Judy's garden i think

OpenStudy (id21):

If you need a proof: x-kim's garden width y-kim's garden length \[x{\times} y\] is Kim's garden area 2x-Judy's garden width y- Judy's garden length 2xy - Judy's garden area. Kim's garden perimeter = 2(x+y)= 2x+2y Judy's garden perimeter = 2(2x+y)=4x+2y "Kim’s garden has the same perimeter as Judy’s garden"- false "The length of Kim’s garden is double the length of Judy’s garden"-false "Kim’s garden has the same length as Judy’s garden"- true "The length of Kim’s garden is four times the length of Judy’s garden"- false